OpenStudy (anonymous):

Add. 0.15x^4 + 0.10x^3 - 0.9x^2 - 0.01x^3 + 0.01^2 + x 1.25x^4 + 0.11x^2 + 0.01 0.27x^3 + 0.99 -0.35x^4 + 15x^2 -0.03

Directrix (directrix):

I cannot follow what we are to add.

OpenStudy (anonymous):

this is how the problem was set up and it just says to add

Directrix (directrix):

Okay, the like terms have already been grouped. I should have noticed. Sorry.

OpenStudy (anonymous):

thats ok, can you help me out please?

Directrix (directrix):

Yes, add the three terms with x^4 in them. They are like terms. Post what you get here. We'll do this in parts.

OpenStudy (anonymous):

0.15x^4+1.25x^4+-0.35x^4=1.75x^4

Directrix (directrix):

0.15x^4 + 1.25x^4 - .35x^4 = 1.05 x^4, I think. Did you subtract the -.35?

OpenStudy (anonymous):

no i added it

OpenStudy (anonymous):

0.10x^3+0.01x^3+0.27x^3=0.38x^3

Directrix (directrix):

Let's look at the coefficients: .15 +1.25 -.35 = 1.40 - .35 = 1.05. 1.05 x^4. We need to agree on the x^4 terms before we move to the x^3. Please check your work with the x^4 terms. Thanks.

OpenStudy (anonymous):

so you add those together and then subtract the .35?

Directrix (directrix):

Yes

OpenStudy (anonymous):

ok got it

Directrix (directrix):

0.10x^3 - 0.01x^3 + 0.27x^3 = .36 x^3 . I will check my answer for the cubes. Please check yours.

OpenStudy (anonymous):

ok i added the 0.01x^3 instead of subtracting thats why my answer was 0.38^3 instead of 0.36^3

OpenStudy (anonymous):

im following you now

Directrix (directrix):

Hey, I'm counting on you to check my work, too. These decimals are always tedious. Okay, next up: the squares. Here's what I got: - 0.9x^2 + 0.01x^2 + + 0.11x^2 = -.78 x^2 And, you?

OpenStudy (anonymous):

you left out the 0.15x^2

OpenStudy (anonymous):

i got -1.17x^2

Directrix (directrix):

Good eye. I missed it. - 0.9x^2 + 0.01x^2 + + 0.11x^2 + 15x^2 = 14.22x^2. Hey, that is a 15x^2 and not .15x^2, right?

OpenStudy (anonymous):

its suppose to be 0.15x^2

OpenStudy (anonymous):

sorry

Directrix (directrix):

- 0.9x^2 + 0.01x^2 + + 0.11x^2 + .15x^2 = -.63x^2

OpenStudy (anonymous):

i dont know what i did wrong on this one but i didnt get that

OpenStudy (anonymous):

i think i added a extra 0 to 0.9x^2

OpenStudy (anonymous):

yes thats where i messed up

Directrix (directrix):

-.9 + .01 = -.89 .11 + .15= .26 ------- -.89 +.26 = -.63 for the x^3 coefficient

OpenStudy (anonymous):

yea i added a 0 and i shouldnt have

OpenStudy (anonymous):

0.01+0.99-0.03=.97

Directrix (directrix):

Okay, is there only one x term? If so, then, 1x for that and we need to do the constants term sum.

OpenStudy (anonymous):

yea it is just x

Directrix (directrix):

0.01 + 0.99 -0.03= -.63

OpenStudy (anonymous):

how do we do the constants term sum?

Directrix (directrix):

These are the constants. Constants don't have an x in them. 0.01+0.99-0.03=.97

OpenStudy (anonymous):

ok gotcha

Directrix (directrix):

My constant sum is not right. Let me check.

OpenStudy (anonymous):

ok

Directrix (directrix):

0.01 + 0.99 -0.03 = .97 Okay, let's put it all together.

OpenStudy (anonymous):

1.05x^4+0.36x^3+-0.63x^2+.97=

Directrix (directrix):

0.15x^4 + 1.25x^4 - .35x^4 +0.10x^3 - 0.01x^3 + 0.27x^3 - 0.9x^2 + 0.01x^2 + + 0.11x^2 + .15x^2 + x + + 0.01 + 0.99 -0.03= 1.05 x^4 +.36x^3 - .63x^2 +x + .97. That's all we can do because these are not like terms and can't be further simplified.

Directrix (directrix):

Did you forget the x?

OpenStudy (anonymous):

oh ok i thought i had to add them all together to get my final answer

OpenStudy (anonymous):

yes i did

Directrix (directrix):

The x is part of the sum. We can't add all of this string together because the terms all have different "last names." By that I mean, x^4, x^3, and so forth. Note that the ones we added to simplify were all of the same last name.

Directrix (directrix):

This is it for the answer: 1.05 x^4 +.36x^3 - .63x^2 + x + .97.

OpenStudy (anonymous):

ok im following you and you are a great tutor. Are you on here all the time because I am terrible at math and you have helped me understand this.

Directrix (directrix):

I try to be on here often. If we "fan" each other, we can tell when we are on by the green dot.

OpenStudy (anonymous):

ok because I could really use your help

Directrix (directrix):

See the attached file. Five of my fans are online now. But you have to fan me if you want to see if I am online.

Directrix (directrix):

Click on the green Directrix and then you'll see a tab to click for "Become a Fan."

OpenStudy (anonymous):

ok thanks